OpenStudy (anonymous):

Simplify the expression 4squareroot256z^16. Assume that all variables are positive. This is the rational and radical functions unit.

OpenStudy (jdoe0001):

\( 4*\sqrt{256}*\sqrt{z^{16}} \implies 4*\sqrt{16^2}*\sqrt{z^{16}} \) and thus \(\sqrt{16^2}=16\) so :)

OpenStudy (anonymous):

Would I then have 4*16*z? By the way THANK YOU!!!

OpenStudy (jdoe0001):

:D

OpenStudy (jdoe0001):

well, one more step

OpenStudy (anonymous):

So my answer would be 64z, correct?

OpenStudy (jdoe0001):

keep in mind that \( z^{16} \implies z^{8*2} \implies z^8*z^2\)

OpenStudy (jdoe0001):

not sure if that made sense... wait a bit

OpenStudy (anonymous):

It made sense. I'm still trying to wrap my head around it though.

OpenStudy (jdoe0001):

\(z^{16} \implies z^{8*8} \implies \pmatrix{z^8}^2\), that, thus you can also get \(z^8\) out

OpenStudy (anonymous):

OH I see where I went wrong. I saw z^16 as z^2, and that's how I got just z when I got the square root...

OpenStudy (jdoe0001):

$$ 4*\sqrt{256}*\sqrt{z^{16}} \implies 4*\sqrt{16^2}*\sqrt{z^{16}}\\ z^{16} \implies z^{8*8} \implies \pmatrix{z^8}^2\\ \text{then }\\ \\ 4\sqrt{16^2\pmatrix{z^8}^2} $$

OpenStudy (anonymous):

so \[\sqrt{z^16}\] completely simplified would be  ? Also sorry my typing isn't very explanatory, I am just joining this site and am not quite sure how to make my equations clear